On 15 July 2016 Sammy Corporation’s Accounts Receivable had a balance of $2300 and the Allowance for Doubtful Debts had a balance of ($220). A specific account of $80 was written off on 16 July 2016. If no other relevant transactions had taken place, what is the amount of net receivables after the write-off?

The answer should be $2080 why?
Answer #1

Accounts Receivable after write off = $2300 - 80 = $2220
Allowance for Doubtful Debts = $220 - 80 = $140

In write off, Accounts Receivable is credited and allowance is debited

Net receivables after the write-off = $2220-140 = $2080